I need help on try to figure out how to evaluate an expression 48 over 2x + 4

note:x=1 x=2 x=3 x=4

Answer by edjones(8007)

48/(2x+4) when x=4
48/((2*4)+4)
=48/(8+4)
=48/12
=4
.
Now you should be able to do the others.
.
Ed